OpenStudy (anonymous):

Solve the exponential equations. (a) 25^(x+1)+5^(2x)=26 (b) 3(3^x)-3^(x-1)=216

myininaya (myininaya):

\[5^{2(x+1)}+5^{2x}=25+1\] \[5^{2(x+1)}+5^{2x}=5^2+5^0\] so looking at both sides both sides would be the same if x=0

myininaya (myininaya):

so x=0 is a solution to the first one

OpenStudy (anonymous):

So that's just something you have to eyeball?

OpenStudy (anonymous):

A similar approach\[25*25^x+25^x=26\]\[25^x(25+1)=26\]\[25^x=1\]\[x=0\]

OpenStudy (anonymous):

ohhhh, clever

myininaya (myininaya):

you can use what mandolino did to other question

myininaya (myininaya):

\[3(3^x)-3^x3^{-1}=216=>3^x(3-3^{-1})=216\]

OpenStudy (anonymous):

The solution of (a) is x=1 (b) is x=4 I KNOW THE ANSWERS BUT I DON'T KNOW HOW TO DO IT.

OpenStudy (anonymous):

\[3*3^x-3^x/3=216\]\[3^x(3-1/3)=216\]\[3^x(8/3)=216\]\[3^x=81\]\[3^x=3^4\]\[x=4\]

myininaya (myininaya):

\[3^x(3-\frac{1}{3})=216=> 3^{x}(\frac{9-1}{3})=216=> 3^x=216 \cdot \frac{3}{9-1}\] \[3^{x}=216 \cdot \frac{3}{8}=>3^{x}=27(3)=>3^x=3^33=>3^x=3^4=>x=4\]

OpenStudy (anonymous):

Same same

myininaya (myininaya):

tissue the answer to a is x=0

OpenStudy (anonymous):

I' m sorry that I typed the QUESTION (a) wrongly. (a) should be 25^(x-1)+5^(2x)=26 AND THE ANSWER IS x=1

OpenStudy (anonymous):

Do you need to see the solution, or are the example enough?

OpenStudy (anonymous):

The example is ENOUGH. THX:)

OpenStudy (anonymous):

Actually this is earlier than your typo equation (a)! \[25^x(1/25+1)=26\]\[25^x(26/25)=26\]\[25^x=26(25/26)\]\[25^x=25^1\]\[x=1\]

OpenStudy (anonymous):

it is very kind of you to help me.THANK YOU SO MUCH!

OpenStudy (anonymous):

I wrote "earlier" I meant "easier"!! I must be getting sleepy!

OpenStudy (anonymous):

25^(x-1)+5^(2x)=26 AND THE ANSWER IS x=1

OpenStudy (anonymous):

5^2(x-1) +5^2x=5^2 +5^0 2(x-1)=2 and 2x=0 power coefficients x-1=1 x=2 ans and 2x=0 x=0 ans